### v1

# Analysis Of A Long Memory Circular Convolution Model

Abstract.
A stochastic model, the product of a circulant matrix and a random normal vector, is shown to produce an evolutive long memory time series with a power law spectral density. The distribution of the time series, a beta location scale family of distributions, provides a connection to the unit centered spherical distribution and directional statistics. The eigenanalysis of the deterministic circulant matrix is shown to provide estimates of the discrete Fourier spectral trend, the intrinsic dimension, the probability density shape parameter of the resulting time series, the condition number of the matrix, and a principle component analysis. Examples of the R code, used as the exploratory element of the work, are given as constructive elements of the paper. The R code may be copied, pasted into an R editor, and explored.
1. Introduction
Peter M. Robinson [1] stated that it is commonly understood that the discrete Fourier spectrum of a long memory time series has a pole (is divergent) at frequency zero. Rafal Weron [2] equates long memory to $1/f$ noise, meaning the power spectral density has a $|frequency|^{-\beta}$ power law curve, which, for $\beta>0$, is also divergent at frequency zero.
This paper adopts the $|frequency|^{-\beta}$ definition of a long memory time series, where $\beta$, the slope of the log log graph of the power spectral density, can assume any value between zero and ten. However, as all sets of quantities described in this paper are bounded in both length and magnitude, the lowest discrete Fourier spectral frequency may approach zero as the length of the model increases, but never becomes zero. The frequency approach to zero, for time series with $\beta>0$, produces an increase in the power spectral density, the distribution of variance with respect to frequency. The result is an increase of variance as the length of the time series increases. The long memory time series is non stationary with respect to variance.
This paper’s model uses a circulant convolution [3] to generate the long memory time series. The eigenvectors of the circulant matrix that generates the convolution are functions of the discrete Fourier transform basis vectors and the eigenvalues of the matrix are the coefficients of the discrete Fourier transform. It will be shown that both the intrinsic dimension, and the probability density function, of the long memory time series are functions of the number of significant low frequency spectral components of the Fourier transform of the time series.
2. The R Language Model that Produces Long Memory Sample Realizations
# long memory time series
# load package named "magic" to produce circulant matrix
beta = 7
n = 500
seed = 5
rn = 0
q = seq(from = -1/2, to = -1/n, by = 1/n)
r = 1/(2*n)
s = -rev(q)
frequency = c(q,r,s)
rn = length(frequency)
density = (abs (frequency))^(- beta/2)
b = Mod((rn^(-1/2)) * fft(density, inverse =TRUE))
# set.seed( seed)
epsilon = rnorm(rn, mean = 0, sd = 1)
circ = (rn^(-1/2)) * circulant (b)
longmem =  circ  %*%  epsilon
plot(longmem, type="l", lwd =2)
Figure 1. Long Memory Time Series From Circulant Matrix Model
2.1. The Model Components Explained
Load a required R language package, and variables are set by user:
  • Load the R language package ”magic” to allow production of circulant matrix.
  • The variable ”beta”, in the interval zero to ten , the slope of the log log plot of the power spectral density of the time series, is set by user.
  • The integer variable ”n”, the reciprocal of the increment of the spectral range of -1/2…-1/n, 1/2n,1/n…1/2 is set by user. This is the Nyquist spectral range of a time series sampled at unit time intervals.
  • If set.seed( ) is to be used, set seed to an integer value.
  • The measured lengths of the spectral range and resulting time series, ”rn”, is initialized to 0 . If ”n” is odd ”rn” will be equal to ”n”. if ”n” is even ”rn” will be ”n”+1. ”rn” will be set later by the program.
The construction of the variable ”frequency” which is the spectral range.
  • Variable ”q” is the vector of negative frequencies.
  • Variable ”r” is the center frequency $1/2n$.
  • Variable ”s” is the vector of positive frequencies.
  • The variable”frequency” is the concatenation of ”q”,”r”, and ”s”, the Nyquist spectral range.
Deriving the first row of the circulant matrix:
  • Variable ”rn” is set to the measured length of ”frequency”.
  • The variable ”density” is set to $|frequency|^{-\beta/2}$. ”density” is the sequence of the moduli of Fourier coefficients of the time series.
  • The variable ”b” is the moduli of the inverse unitary discrete Fourier transform of ”density”.
The normal random vector generated , the circulant matrix produced, and the matrix vector product produces the long memory time series.
  • Set.seed( ) may be uncommented by removing the hash sign # to allow multiple extended sequences of ”longmem”.
[…]
  • The circulant matrix ”circ” is produced from ”b” using the ”circulant” command from the package ”magic”, and normalized by one over the square root of ”rn”.
  • The matrix vector product of ”circ” and ”epsilon” completes the circular convolution that produces the long memory time series ”longmem”. A plot of ”longmem” is created.
3. The Symmetric Circulant Matrix
The first row of the circulant matrix ”circ”, produced by the model parameters ”rn” = 7 and ”beta” = 3 is:
(12.510, 8.253, 6.140, 5.543, 5.543, 6.140, 8.253)
The sequence, after the left most element, is a palindromic sequence of numbers, such as 1,2,3,3,2,1 or 0,0,0,0 or 1,1,1,1. The second row of the circulant matrix is the first row circularly shifted to the right, with the last element in the row moved to the first position. The next rows follow the same pattern of a circular shift of the row above. An example of a symmetric circulant matrix of letters is :
$\left[\begin{matrix}A&B&C&B\\ B&A&B&C\\ C&B&A&B\\ B&C&B&A\\ \end{matrix}\right]$
The matrix is a symmetric matrix, symmetrical about its top left to lower right diagonal.The matrix is self adjoint, equal to its transpose. The first row of the matrix is identical to the first column, the second row is identical to the second column, and so on. The square symmetric positive semi-definite circulant matrix used to model long memory time series in this paper has an odd number of columns and rows. The matrix is multiplied by $rn^{-0.5}$ to produce an unitary transformation.
3.1. The Top Row of The Circulant Matrix
”beta” top row ”circ” ”density”
0 1, 0, 0, 0, 0 1, 1, 1, 1, 1
7 664, 637, 612, 612, 637 11.3, 67.6, 3162, 67.6, 11.3
Table 1. The Response of Fourier Transform Pair ”Circ” top row, and ”Density” to ”Beta”
Table 1 is a chart of the top row of the normalized matrix ”circ” and the variable ”density” in response to two values of ”beta”. The length of the model is five. The top row of ”circ” and the variable”density” are a scaled Fourier transform pair, giving insight to the response of the model over the range of ”beta”.
When ”beta” is zero, the the frequency domain variable ”density” is a vector of numbers raised to the zero power, which is a vector of ones. The inverse Fourier transform of ”density” is the time domain top row of the matrix ”circ”, a vector of one followed by zeros. The matrix ”circ”, after the circular permutation of the top row, becomes an identity matrix and the resulting time series after circular convolution, is the vector ”epsilon”.
When ”beta” is seven, the frequency domain variable ”density” is a vector with large value at nearly zero frequency and smaller values symmetrically located on either side. The time domain variable, the top row of ”circ” is a vector of nearly similar values. The matrix ”circ”, after the circular permutation of the top row, becomes a nearly singular matrix with large condition number.
4. Equivalence Class of Location and Scale transforms of a Distribution
Michel Loeve [4] stated in ”Probability Theory” that the type or law of a distribution function is independent of location and scale. Expressed differently, probability distributions differing only by location and or scale transformations are an equivalence class. This definition of distribution type of a long memory time series is used in this paper. Some items to remember are:
  • The dot product of two vectors A and B is equal to the product of the $L^{2}$ norms of A and B and the cosine of the angle between A and B. The dot product of A and B = $\parallel A\parallel\parallel B\parallel cos\theta$.
  • Each component of vector ”longmem” is the dot product of a row of the matrix ”circ” and the vector ”epsilon”.
  • The norms of each row of ”circ” are equal. Each row of ”circ” is composed of a permutation of the same numbers.
[…]
Dividing the vector ”longmem” by the product of the two $L^{2}$ norms in equation (1) yields an equivalent location scale distribution.
It is interesting to note that the probability distribution of the vector of cosines of angles between the unit magnitude rows of ”circ” and the unit magnitude vector ”epsilon” is a directional statistic.
It will be shown in section 6 that the probability distribution of the vector of cosines has a general beta distribution as defined by NIST ( National Institute of Standards and Technology) [5] on their online EDA Handbook ( John Tukey’s Exploratory Data Analysis). NIST defines the lower limit of the values of the general beta distribution, ”a”, as a location parameter, ”b” as the upper limit, and the range, ”b - a” , as the scale parameter. In order to arrive at a standard beta distribution, defined as having support on the interval [0,1], each component of the cosine vector in equation (1) has location parameter ”a” subtracted from it to produce a lower limit of zero, then the scale parameter ”b-a” divides the result of the subtraction to produce an upper limit of one. The result is shown in equation (2).
(2) $\left[\begin{matrix}(cos\theta_{1}-a)/(b-a)\\ \vdots\\ (cos\theta_{n}-a)/(b-a)\end{matrix}\right]$
5. Histograms of Long Memory Time Series at different values of ”beta”
The vector of equation (2) is a small sample of an autocorrelated time series, so in the following R program, many standardized vectors of equation (2) will be concatenated to produce an adequate representation of the probability density function.
5.1. The R Language Program to Produce A Histogram
#distribution of a long memory time series
# load R package "magic"
beta = 2.3
n = 200
result = 0
q = seq(from = -1/2, to = -1/n, by = 1/n)
r = 1/(2*n)
s = - rev(q)
frequency = c(q,r,s)
rn = length(frequency)
density = (abs(frequency))^(-beta/2)
b = Mod((rn^(-1/2)) * fft(density, inverse = TRUE))
circ = (rn^(-1/2)) * circulant (b)
for(i in 1:200) {
epsilon = rnorm(rn, mean = 0, sd = 1)
norm_epsilon = (sum(epsilon^2))^0.5
norm_row = (sum(circ[1, ]^2))^.5
longmem = circ %*% epsilon
cosvec = longmem/ (norm_epsilon * norm_row)
standard = (cosvec - min(cosvec))/(max(cosvec) - min(cosvec))
result = c(result, standard)  }
result = setdiff(result, c(0,1))
hist(result, freq = FALSE, breaks = 100)
5.2. The Histogram Program Explained
Much of this code is identical to the code in section 2.1 . The explanation will begin at the for loop staring with ”for(i in 1:200)”.
  • The for loop ”for(i in 1:200)” starts a for loop from 1 to 200, indexed by i. For each pass through the loop a new vector ”epsilon ” is created. The next two line calculate the vector norms of ”epsilon” and the top row of the circulant matrix ”circ”, these are called ”norm epsilon” and ”norm row”.
  • The time series ”longmem” is created from the matrix product of ”circ” and the vector ”epsilon”. On the next line of code, the vector of cosines ”cosvec” is created by dividing ”longmem” by the product of ” norm epsilon” and ”norm row”.
  • ”cosvec” is standardized as in equation 2 to produce the vector ”standard”.
  • The 200 different vectors ”standard” produced by the for loop are concatenated as the vector ”result”. The vector ”result” then has the excess ones and zeros removed by the setdiff(result, c(1,0) command.
  • The histogram, an empirical density function, is produced with 50 bins from ”result”.
Figure 2. Four Histograms Of Time Series
6. The General Beta Distribution Of Long Memory Time Series
The histograms of figure 2 are generated from the paper’s model of long memory time series with different values of the slope of the power spectral density ”beta”. They represent four standardized density functions in the beta family of distributions. The top left is the truncated normal, the top right is the Wigner semicircle, the bottom left is the uniform, and the bottom right is the arcsine distribution. These distributions are not unique to the model used in this paper, the distributions can be generated by the R language ARFIMA model of Granger and Joyeux as well as the R language model of Timmer and Konig’s 1995 model TK95. Details can be found in an earlier paper on arXiv by the author of this paper.
6.1. The Ratio Of Variance Divided By The Range Squared
The variance of the general beta distribution is given by NIST as:
(3) $\sigma^{2}=(b-a)^{2}\frac{\alpha\beta}{(\alpha+\beta)^{2}(\alpha+\beta+1)}$
Where, a is the lower limit of the range, b is the upper limit of the range and $(a-b)^{2}$ is the range squared. The variables, $\alpha$ and $\beta$, are the two shape parameters of the beta distribution. Fortunately for this paper, the beta distributions in this paper are symmetric, and $\alpha=\beta$, so all references to $\beta$ are replaced by $\alpha$. Dividing both sides of equation (3) by the range squared, replacing $\beta$ with $\alpha$ and simplifying yields:
(4) $\frac{\sigma^{2}}{(b-a)^{2}}=\frac{1}{8\alpha+4}$
A property of the left side of equation 4 is scale invariance. Apostal [6] states that al homogeneous function $f$ of degree $p$ is a function of a sequence $x$ and a scalar $t$ such that :
(5) $f(t(x))=t^{p}f(x)$
Variance and the range squared are both homogeneous functions of degree two, so any scale transformation of the the random sequence $x$ cancels in the ratio. Variance and range are also location invariant. The result is that the variance to range squared ratio, easily measured from finite long memory sequences , can be used to determine the beta shape parameter $\alpha$.
Density Variance / Range Squared Shape Parameter Alpha
Normal Approaching 0 Approaching Infinity
Wigner 1/16 3/2
Uniform 1/12 1
Arcsine 1/8 1/2
Bernoulli 1/4 0
Table 2. Table of Variance Divided By Range Squared, and Beta Distribution Shape Parameter
6.2. Popoviciu’s Inequality of Variance
Tiberiu Popoviciu’s variance inequality on bounded probability distributions [7], with (b - a) equal to the range of the distribution, states:
(6) $\sigma^{2}\leqq\frac{(b-a)^{2}}{4}$
with only the Bernoulli distribution having equality with 1/4.
A proof of Popoviciu’s theorem as applied to symmetric distributions follows.
Note that a symmetric Bernoulli distribution with probability 1/2, has a variance greater than any other bounded distribution with the same range (b - a). The distance from the extreme ends of the Bernoulli distribution, where the entire mass of the random variable values are found, to the mean is $x-\mu$, eqaul to plus or minus (b - a)/2 . The ratio of variance to squared range of the Bernoulli distribution with probability 1/2 is:
(7) $\sigma^{2}$ $=1/2\left[\left(\frac{(b-a)}{2}\right)^{2}+\left(\frac{(a-b)}{2}\right)^{2}\right]=\frac{(b-a)^{2}}{4}$
$\frac{\sigma^{2}}{range^{2}}$ $=\frac{(b-a)^{2}}{4}\frac{1}{(b-a)^{2}}=1/4$
The ratio of variance to range squared of any bounded distribution is $\leqq 1/4$.
6.3. The Unit Centered Sphere and Cosine Similarity
The uniformly distributed unit vector variables ”epsilon” in the long memory histogram model can be represented as a $S^{d-1}$ spherical surface, with d $\in\mathbb{N}$. The unit row vectors of the circulant matrix are also members of the unit hypersphere. What determines the particular beta distribution is the intrinsic dimension d of the dot product of the unit rows of the circulant matrix and the unit normal vectors ”epsilon”. Four references derive the beta distribution on the unit d dimensional sphere. Djalil Chafai’s two papers [8] [9] use the Funk Hecke formula to derive the beta density function. William A Huber [10] uses a geometric construction on the $S^{d-1}$ sphere to arrive at the same density function. Tony Cai and Tiefeng Jiang [11] derive the density function from the distribution of Pearson correlation coefficients. The density function given by all four references is the following expression:
(8) $f_{t}$ $=\frac{\Gamma(\frac{d}{2})}{\pi^{1/2}\Gamma(\frac{d-1}{2})}(1-t^{2})^{\frac{d-3}{2}}$
$f_{t}$ $\propto(1-t^{2})^{\frac{d-3}{2}}$
The random variable t in equation (8), created by the many dot product cosines, has a support of [-1,1]. An affine transformation of variable t to variable u = (t+1)/2 will have the standard beta distribution support of [0,1]. Below is a template of the linear transformation of the density function of $f_{t}$ to $f_{u}$.
(9) $u$ $=at+b$
$t$ $=(u-b)/a$
$f_{u}$ $=1/|a|\ f_{t}\left[\frac{u-b}{a}\right]$
Insert the values of a = b = 1/2 in the second line of equation (9), where t = 2u - 1 and - t = 1 - 2u. Substituting 1-2u for -t in equation (8) and simplifying the expression yield a beta distribution with support on the interval [0,1].
(10) $t$ $=(u-b)/a,\ a=b=0.5$
$t$ $=2u-1$
$-t$ $=1-2u$
$f_{t}$ $\propto(1-t^{2})^{\frac{d-3}{2}}$
$f_{u}$ $\propto 2\left[1-(1-2u)^{2}\right]^{\frac{d-3}{2}}$
$f_{u}$ $\propto 2\left[1-(4u^{2}-4u+1)\right]^{\frac{d-3}{2}}$
$f_{u}$ $\propto 2\left[4(u-u^{2})\right]^{\frac{d-3}{2}}$
$f_{u}$ $\propto 2^{D-2}\left[u-u^{2}\right]^{\frac{d-3}{2}}$
$f_{u}$ $\propto\left[u-u^{2}\right]^{\frac{d-3}{2}}$
$f_{u}$ $\propto u^{\frac{d-1}{2}-1}(1-u)^{\frac{d-1}{2}-1},\ given\frac{d-3}{2}=\frac{d-1}{2}-1$
$f_{u}$ $\propto u^{\alpha-1}(1-u)^{\alpha-1},with\ \alpha=\frac{d-1}{2}$
The last line of equation (10) demonstrates that $u=(t+1)/2$ has a symmetric beta distribution with shape parameter $\alpha=(d-1)/2$. As $\alpha$ is a continuous variable it is possible to speak about non integer values of intrinsic dimension d given $\alpha=\frac{d-1}{2}$.
The four references state that when d = 2 the beta density is an arcsine, d = 3 is an uniform, and when d = 4 the density function is a Wigner semi-circle. It is apparent that the extrinsic dimensions involved in the histogram model given in section (5) are much greater than d = 2,3, or 4. It will be shown, in the following section, that the intrinsic dimension of the rows of the symmetric circulant matrix match the values given by the four references.
7. Statistical Estimates Derived from the Eigenanalysis of the Symmetric Circulant Matrix
A single autocorrelated long memory time series has a limited amount of statistical information. Sampling many independently produced, equal length, and standardized time series is one solution to derive statistical trends. The histograms of the probability density functions produced by the program in section 5.1 were generated from 200 samples of time series sequences. Another approach is to derive the ratio of variance to squared range of many standardized time series, as demonstrated in section 6.1. Both variance and range are easily measured from the concatenation of many time series, and the ratio of variance to squared range determines the beta parameter $\alpha$, and intrinsic dimension d.
Another solution is to estimate the statistical trends directly from the eigenanalysis of the symmetric circulant matrix used to generate the time series. The eigenvectors are the orthogonal discrete Fourier basis vectors, and the eigenvalues are the coefficients of the discrete Fourier transform.The eigenvalues of a $rn\times rn$ matrix are indexed with respect to magnitude, $\lambda_{1}$ being the largest, and $\lambda_{rn}$ being the smallest. Squaring the moduli of Fourier coefficients produces an estimate of the power spectral density. The mean integral of the power spectral density is an estimate of the variance. It will be shown that trend estimates of the beta parameter $\alpha$, intrinsic dimension d, range, and the condition number of the circulant matrix can be produced from the eigenvalues.
7.1. Estimates of Intrinsic Dimension, Beta Distribution Shape Parameter, and Principle Components
Ipsen and Saibaba [12] define intrinsic dimension of a matrix A as:
[…]
Where trace(A) is the sum of the main upper left to lower right diagonal elements of matrix A, or equally the sum of the eigenvalues of matrix A. The denominator of equation(11) is the spectral norm of matrix A, the largest eigenvalue, $\lambda_{1}$. Equation (11) is a function of the eigenvalues of matrix ”circ”.
The value of intrinsic dimension is context sensitive [13], and for this paper equation(11) is a good first approximation. A better estimate for intrinsic dimension d, between 2 to 4, and $\alpha$, between 0.5 and 1.5, is equation(12), an empirically derived two step estimate.
[…]
$d$ $\cong E-\frac{E-3}{4.2}$
$\alpha$ $\cong\frac{d-1}{2}$
Figure 3. Log Log Plots of Eigenvalues
Figure 3 are two log log plots of eigenvalues for a normal distribution beta = 0.001, and an arcsine distribution, beta = 10. Replacing trace(A) in equation (11) with the sum of eigenvalues, and $A||_{1}$ by the largest eigenvalue $\lambda_{1}$ yields:
(13) $\frac{\lambda_{1}+\sum_{i=2}^{rn}\lambda_{i}}{\lambda_{1}}$
Equation (13) while imprecise, is useful to observe trends at the extremes of the range of ”beta”. When beta approaches zero as in the left plot of figure 3, equation(13) approaches rn, meaning all Fourier moduli are approximately equal and the distribution will approximate a normal distribution if rn is large enough. When beta is ten, the right plot of figure 3, equation (13) approaches unity, and the sum of $\lambda_{1}$ and $\lambda_{2}$ dominate the sum of all eigenvalues.The sum of $\lambda_{1}$ plus $\lambda_{2}$, are 96.8 percent of the sum of all eigenvalues, and they are the principle components of the time series. A time series with a beta of ten has an arcsine distribution, the distribution of single frequency sinusoid. The bottom right plot of figure 1, with beta = 10, is seen to be nearly sinusoidal. The plots in figure 3 represent the Fourier spectral trends of the long memory time series. The Fourier moduli of the time series are randomized by the product of ”circ” with ”epsilon”, but on average, the spectral trend is an accurate statistic with a slope of $\beta/2$. The log log plot of the squared eigenvalues has a slope equal to $\beta$, the trend of the power spectral density.
7.2. Eigenvalue Estimate of The Variance of Long Memory Time Series
Parseval’s theorem, when applied to a unitary discrete Fourier transform, with x the time domain variable, and X the Fourier frequency domain variable, can be expressed as:
(14) $\sum_{j=1}^{n}|x_{j}|^{2}=\sum_{k=1}^{n}|X_{k}|^{2}$
In order to estimate the variance, it is necessary to remove the coefficient of the constant offset, zero frequency Fourier term $X_{1}$. The sum of Fourier trigonometric series without the zero frequency term is zero mean. Translating equation (14) to an estimate of variance using the eigenvalues of ”circ” produces the following equation:
(15) $var(x)\cong\frac{1}{rn-1}\sum_{k=2}^{rn}\lambda_{k}^{2}$
Given the estimates of $\alpha$ and variance, the range of the general beta distribution may be estimated. If the variance of ”epsilon” is different than 1 then the right side of equation (15) needs to be multiplied by the new variance of ”epsilon”.
7.3. The Condition Number of The Symmetric Matrix ”Circ” and Geometric Collapse
A symmetric matrix has a condition number $\kappa$ equal to the maximum eigenvalue divided by the minimum eigenvalue [14].
(16) $\kappa=\frac{\lambda_{1}}{\lambda_{n}}$
The matrix ”circ”, when ”beta” = 0, has a distribution that is normal, and $\kappa$ = 1. When ”beta” = 2.2, the distribution is a Wigner semi-circle, and $\kappa=3.4\times 10^{2}$. When ”beta” = 10, the distribution is an arcsine, and $\kappa=3.2\times 10^{11}$.
The reduction of intrinsic dimension, increase in serial correlation of the time series, decreasing significance of smaller eigenvalues, and approach to singularity of the matrix ”circ”, follows the increase of the condition number $\kappa$. This is called the geometric collapse of the matrix. The intrinsic dimension of the long memory time series is the number of significant Fourier terms. Figure 1 illustrates the gradual reduction of high frequency spectral components as ”beta” increases.
$\beta$ Statistic Eigenvalue Mean Measured Measured
Estimated Statistic Coefficient of
Statistic Variation
2.2 d 4.17 4.22 0.22
3.0 d 2.93 3.21 0.23
10 d 2.05 2.00 0.03
2.2 $\alpha$ 1.58 1.61 0.29
3.0 $\alpha$ 0.97 1.10 0.34
10 $\alpha$ 0.52 0.50 0.05
2.2 variance 1.72 E 3 1.67 E 3 0.71
3.0 variance 9.62 E 4 8.80 E 4 0.88
10 variance 1.03 E 21 1.03 E 21 0.96
Table 3. Statistics Estimated from Eigenvalues, and Measured from 500 Long Memory Sample Sequences of Length 200, using the R code in subsection 7.4
7.4. The R language Program to Evaluate the Estimates of Intrinsic Dimension, Beta Shape Parameter, and Variance Derived from the Eigenvalues of ”Circ”
#Estimates of d, alpha, and variance derived from eigenvalues
#Load package "magic" to produce circulant matrix
beta = 2.2
n = 200
alpha = 0
d = 0
v = 0
q = seq(from = -1/2, to = -1/n, by = 1/n)
r = 1/(2*n)
s = -rev(q)
frequency = c(q,r,s)
rn = length(frequency)
density = (abs(frequency))^(-beta/2)
b = Mod((rn^(-1/2))*fft(density, inverse = TRUE))
circ = (rn^(-1/2)) * circulant(b)
e = eigen(circ)
val = e\$values
# ESTIMATION OF D AND ALPHA FROM EIGENVALUES
E = (sum(val[1:rn])/((2^(1/2))*val[1]))+1
d_est = E - (E-3)/4.2
alpha_est = (d_est - 1)/2
#ESTIMATION OF VARIANCE FROM EIGENVALUES
var_est = (1/(rn-1)) * sum(val[2:rn]^2)
#LOOP TO MEASURE ACTUAL MEAN AND COEFFICIENT
#OF VARIATION OF D, ALPHA, AND VARIANCE
for(i in 1:500) {
epsilon = rnorm(rn, mean = 0, sd = 1)
longmem = circ%*% epsilon
# 500 MEASURES OF ALPHA, D, AND VARIANCE
alpha[i]= (1/8) * ((max(longmem)-min(longmem))^
2 / var(longmem)) -1/2
d[i] = 2 * alpha[i] +1
v[i] = var(longmem) }
#ESTIMATES OF D, AND ALPHA FROM EIGENVALUES
alpha_est
d_est
formatC(var_est, format = "e")
#MEASURED MEAN, AND COEFFICIENT OF
#VARIATION FROM 500 SAMPLES
mean(alpha)
sd(alpha)/mean(alpha)
mean(d)
sd(d)/mean(d)
formatC(mean(v), format = "e")
sd(v)/mean(v)
8. Acknowledgment
The author found inspiration to pursue this topic and write this paper from the work of Herman Wold’s 1938 book ”A Study In The Analysis Of Stationary Time Series” [15], as well as the ubiquity of 1/f noise in the outputs of astronomical instrumentation.